# HoTPP Benchmark: Are We Good at the Long Horizon Events Forecasting?

### v3

@@
Abstract.
Forecasting multiple future events within a given time horizon is essential for applications in finance, retail, social networks, and healthcare. Marked Temporal Point Processes (MTPP) provide a principled framework to model both the timing and labels of events. However, most existing research focuses on predicting only the next event, leaving long-horizon forecasting largely underexplored. To address this gap, we introduce HoTPP, the first benchmark specifically designed to rigorously evaluate long-horizon predictions. We identify shortcomings in widely used evaluation metrics, propose a theoretically grounded T-mAP metric, present strong statistical baselines, and offer efficient implementations of popular models. Our empirical results demonstrate that modern MTPP approaches often underperform simple statistical baselines. Furthermore, we analyze the diversity of predicted sequences and find that most methods exhibit mode collapse. Finally, we analyze the impact of autoregression and intensity-based losses on prediction quality, and outline promising directions for future research. The HoTPP source code, hyperparameters, and full evaluation results are available on GitHub [^1] https://github.com/ivan-chai/hotpp-benchmark.
copyright: acmlicensedjournalyear: 2025
1. Introduction
@@
Event sequences differ fundamentally from other data types. Unlike tabular data (Wang and Sun, 2022), ESs include timestamps and possess an inherent order. In contrast to time series data (Lim and Zohren, 2021), ESs are characterized by irregular time intervals and additional data fields, such as event types. These differences necessitate the development of specialized models and evaluation practices.
@@
Irregularly spaced timestamps are commonly modeled via Temporal Point Processes (TPP) (Rizoiu et al., 2017). Assigning categorical labels to these events leads to the Marked Temporal Point Process (MTPP) formalism. Since most event-sequence data can be viewed as MTPPs augmented with additional fields (McDermott et al., 2024), accurately modeling event times and labels remains a central challenge in the broader event-stream (ES) domain.
@@
In practice, a common question arises: what events will occur, and when, within a specific time horizon? Forecasting multiple future events presents unique challenges that differ from traditional next-event prediction tasks (Xue et al., 2022; Zhou et al., 2025). For instance, autoregressive event sequence prediction involves applying the model to its own potentially erroneous predictions. However, the challenges of autoregressive prediction in the context of MTPP have not been thoroughly explored. Another difficulty lies in evaluation. Methods like Dynamic Time Warping (DTW) are typically unsuitable for event sequences due to strict ordering constraints (Su and Hua, 2017). Some works have applied Optimal Transport Distance (OTD), a variant of the Wasserstein distance, to compare sequences of predefined lengths (Mei et al., 2019; Xue et al., 2022, 2023). However, the limitations of this metric have not been previously considered.
@@
Figure 1. T-mAP evaluation pipeline. Unlike previous approaches, T-mAP evaluates sequences of variable lengths within the prediction horizon. It further improves performance assessment by analyzing label distributions rather than relying on fixed predictions.
@@
In this work, we provide the first in-depth analysis of models and metrics for long-horizon event forecasting, establishing a rigorous evaluation framework and a baseline for MTPP studies. Our key contributions are summarized as follows:
  1. (1)
@@
    We introduce HoTPP, the first open-source benchmark specifically designed for long-horizon event forecasting. By considering datasets and methods spanning finance, social networks, healthcare, and recommender systems, HoTPP significantly broadens the diversity and scale of data compared to previous benchmarks (Xue et al., 2023; McDermott et al., 2024; Xu, 2018). Moreover, we provide optimized training and inference algorithms to support large-scale studies.
  2. (2)
@@
    We demonstrate that widely used evaluation metrics for MTPPs overlook critical aspects of model performance. In particular, they are sensitive to calibration and incorrectly count false positives and false negatives. To address these issues, we introduce Temporal mean Average Precision (T-mAP), a novel evaluation metric inspired by best practices in computer vision (Lin et al., 2014). Furthermore, we address a theoretical gap in computer vision by proving the correctness of the T-mAP computation algorithm.
  3. (3)
@@ 2. Related Work
Marked Temporal Point Processes. A Marked Temporal Point Process (MTPP) is a stochastic process consisting of a sequence of pairs $(t_{1},l_{1}),(t_{2},l_{2}),\dots$, where $t_{1}<t_{2}<\dots$ represent event times and $l_{i}\in\{1,\dots,L\}$ denote event type labels (Rizoiu et al., 2017). Common approaches to MTPP modeling focus on predicting the next event. A basic solution involves independently predicting the event time and type. A more advanced approach splits the original sequence into subsequences, one for each event type, and independently models each subsequence’s timing. Depending on the form of time-step distribution, the process is called Poisson or Hawkes.
@@
Over the last decade, the focus has shifted toward increasing model flexibility by applying neural architectures. Several works have employed classical RNNs (Du et al., 2016; Xiao et al., 2017; Omi et al., 2019) and transformers (Zuo et al., 2020; Zhang et al., 2020), while others have proposed architectures with continuous time (Mei and Eisner, 2017; Rubanova et al., 2019; Kuleshov et al., 2024). In this work, we evaluate MTPP models with both discrete and continuous time architectures. Unlike previous benchmarks, we also assess simple rule-based baselines and popular methods from related fields, including GPT-like prediction models for event sequences (McDermott et al., 2024; Padhi et al., 2021). For more details on MTPP modeling, refer to Appendix A.
@@
MTPP Evaluation. Previous benchmarks have primarily focused on medical data or traditional MTPP datasets. EventStream-GPT (McDermott et al., 2024) and TemporAI (Saveliev and van der Schaar, 2023) consider only medical data and do not implement methods from the MTPP field, despite their applicability. Early MTPP benchmarks, such as Tick (Bacry et al., 2017) and PyHawkes [^2] https://github.com/slinderman/pyhawkes, implement classical machine learning approaches but exclude modern neural networks. While PoPPy (Xu, 2018) and EasyTPP (Xue et al., 2023) include neural methods, they do not consider rule-based and horizon prediction approaches, such as HYPRO (Xue et al., 2022) and Diffusion (Zhou et al., 2025). Furthermore, PoPPy does not evaluate long-horizon predictions at all. EasyTPP addresses this limitation to some extent by reporting the OTD metric, though it does not provide the corresponding evaluation code.
@@
Previous work also highlighted the limitations of Dynamic Time Warping (DTW) for event sequence evaluation (Su and Hua, 2017), showing that DTW’s strict ordering constraints are impractical and should be avoided. This work proposed an alternative, the Order-preserving Wasserstein Distance (OPW). In our study, we address the limitations of OTD, a variant of the Wasserstein Distance, and propose a new metric called T-mAP. Unlike OPW, T-mAP operates on timestamps rather than event indices and involves two hyperparameters, compared to three in OPW.
Figure 2. Comparison of the OTD and T-mAP metrics. In the example 1.a, OTD’s prefix-based evaluation of three events erroneously produces a false negative, whereas T-mAP in example 1.b compares all events within the specified time horizon for a more accurate assessment. Moreover, OTD evaluates only the label with the highest probability, treating cases 2.a and 2.b as equivalent. By contrast, T-mAP considers the entire distribution, yielding a low score (0.33) for 2.a and a perfect score (1) for 2.b.
3. T-mAP Metric
@@
In this section, we show that widely used evaluation metrics for long-horizon prediction often miscount false positives (FP) and false negatives (FN). We then introduce Temporal mean Average Precision (T-mAP), a novel, theoretically grounded metric that addresses these issues. In addition, T-mAP is invariant to linear calibration of model outputs, simplifying and stabilizing comparisons across different methods. Additional T-mAP experiments on sequences with highly irregular time steps are presented in Appendix H. Long-tail performance measurement is discussed in Appendix I.
3.1. Limitations of the Next-Event and OTD Metrics
@@
Recent studies have advanced the measurement of horizon prediction quality by employing OTD (Mei et al., 2019). This metric is analogous to the edit distance between the predicted event sequence and the ground truth. Suppose there is a predicted sequence $S_{p}=\{(t^{p}_{i},l^{p}_{i})\}_{i=1}^{n_{p}},0<t_{1}^{p}<t_{2}^{p}<\dots$, and a ground truth sequence $S_{gt}=\{(t^{gt}_{j},l^{gt}_{j})\}_{j=1}^{n_{gt}},0<t_{1}^{gt}<t_{2}^{gt}<\dots$. These two sequences form a bipartite graph $\mathcal{G}(S_{p},S_{gt})$. The prediction is connected to the ground truth event if their types are equal. Denote $M(S_{p},S_{gt})$ a set of all possible matches in the graph $\mathcal{G}(S_{p},S_{gt})$. OTD finds the minimum cost among all possible matchings:
(1) $\mathrm{OTD}(S_{p},S_{gt})=\\ \min\limits_{m\in M(S_{p},S_{gt})}\left[\sum\limits_{(i,j)\in m}|t^{p}_{i}-t^{gt}_{j}|+C_{del}U_{p}(m)+C_{ins}U_{gt}(m)\right],$
where $U_{p}(m)$ is the number of unmatched predictions in matching $m$, $U_{gt}(m)$ is the number of unmatched ground truth events, $C_{ins}$ is an insertion cost, and $C_{del}$ is a deletion cost. It is common to take $C_{ins}=C_{del}$ (Mei et al., 2019). It has been proven that OTD is a valid metric, as it is symmetric, equals zero for identical sequences, and satisfies the triangle inequality.
@@
3.2. Temporal mAP: Definition and Computation
In this section, we introduce a novel metric, Temporal mAP (T-mAP), which analyses all errors within a predefined time horizon, explicitly controls ordering, and is invariant to linear calibration. T-mAP is inspired by object detection methods from computer vision (Everingham et al., 2010), as illustrated in Fig. 1. Object detection aims to localize objects within an image and identify their types. In event sequences, we tackle a similar problem but consider the time dimension instead of horizontal and vertical axes. Unlike object detection, where objects have spatial size, each event in an MTTP is a point without a duration. Therefore, we replace the intersection-over-union (IoU) similarity between bounding boxes with the absolute time difference.
@@ 3.2.1. Definition
Figure 4. T-mAP dependency on the $\delta$ parameter. The dashed line indicates the selected value.
3.2.2. Computation
Finding the optimal matching independently for each threshold value $h$ is impractical; thus, we need a more efficient method to evaluate T-mAP. This section shows how to optimize T-mAP computation using an assignment problem solver, like the Jonker-Volgenant algorithm (Jonker and Volgenant, 1988). The resulting complexity of T-mAP computation is $\mathcal{O}(BN^{3})$, where $B$ is the number of evaluated sequences and $N=\max(n_{p},n_{gt})$ is the number of events within the horizon.
@@ 3.2.2. Computation
(3) $\forall h\forall m\in M(\mathcal{G})\exists m_{h}\in M(\mathcal{G}_{h}):m_{h}\subset m.$
According to this theorem, we can compute the matching for the prediction $S^{l}_{gt}$ and subsequently reuse it for all thresholds $h$ and subsequences $S^{l}_{>}(h)$ to construct a precision-recall curve for the entire dataset, as shown in Fig. 3.c. The proof of the theorem, the study of calibration dependency, and the complete algorithm for T-mAP evaluation are provided in Appendix B.
3.2.3. T-mAP Hyperparameters
@@
4. HoTPP Benchmark
@@
The HoTPP benchmark integrates data preprocessing, training, and evaluation in a single toolbox. Unlike previous benchmarks, HoTPP introduces the novel T-mAP metric for long-horizon prediction. HoTPP also differs from prior MTPP benchmarks by including simple rule-based baselines and long-horizon prediction methods that simultaneously predict multiple future events, such as HYPRO (Xue et al., 2022) and Diffusion (Zhou et al., 2025). The HoTPP benchmark is designed to focus on simplicity, extensibility, evaluation stability, reproducibility, and computational efficiency.
@@
Simplicity and Extensibility. The benchmark code is organized into a clear structure, separating the core library, dataset-specific scripts, and configuration files. New methods can be integrated at various levels, including the configuration file, model architecture, loss function, metric, and training module. Core components are reusable through the Hydra configuration library (Yadan, 2019). For example, HoTPP can apply the HYPRO rescoring method to backbone models such as IFTPP, RMTPP, NHP, or ODE. Furthermore, each method’s backbone architecture can be easily switched between RNNs and Transformers. Methods such as IFTPP, RMTPP, and Diffusion are compatible with most HuggingFace models (Wolf et al., 2020).
@@
Evaluation Stability. In the MTPP domain, many datasets contain only a few thousand sequences for training and evaluation. For example, the StackOverflow test set includes 401 sequences, Retweet contains 1956, and Amazon has 1851 sequences for testing. This possesses a unique challenge in the long-horizon forecasting task. Previous long-horizon evaluation pipelines (Xue et al., 2022, 2023) made predictions only at the end of each sequence, resulting in a limited number of measurements and a low evaluation stability. To address this limitation, we evaluate long-horizon predictions at multiple intermediate points, decreasing the variance of evaluation metrics.
Table 1. Datasets statistics
Dataset # Sequences # Events Mean Length Mean Horizon Mean Duration # Classes OTD T-mAP Length Steps Cost $\delta$ Horizon Transactions 50k 43.7M 875 9.0 719 203 5 1 2 7 MIMIC-IV 120k 2.4M 19.7 6.6 503 34 5 2 4 28 Retweet 23k 1.3M 56.4 14.7 1805 3 10 15 30 180 Amazon 9k 403K 43.6 14.8 22.1 16 5 1 2 10 StackOverflow 2k 138K 64.2 12.0 55.3 22 10 1 2 10
@@
Reproducibility. The HoTPP benchmark ensures reproducibility at multiple levels. First, we use the PytorchLightning library (Falcon and The PyTorch Lightning team, 2019) for training with a specified random seed and report multi-seed evaluation results. Second, data preprocessing is carefully designed to ensure datasets are constructed reproducibly. Finally, we specify the environment in a Dockerfile.
Table 2. Evaluation results. The best result is shown in bold. The mean and standard deviation of each metric computed during five runs with different random seeds are reported.
@@
Computational Efficiency. Some methods are particularly slow, especially during autoregressive inference on large datasets. Straightforward evaluation can take several hours on a single Nvidia V100 GPU. We optimized the training and inference pipelines in two ways to accelerate computation. First, we implemented an efficient RNN that reuses computations during parallel autoregressive inference from multiple starting positions. Second, we developed highly optimized versions of the thinning algorithm used for sampling in NHP and continuous-time neural architectures (NHP and ODE) (Rizoiu et al., 2017), achieving up to a 4x performance improvement compared to the official implementations. These optimizations allowed us to conduct the first large-scale evaluations of algorithms such as NHP, AttNHP, ODE, and HYPRO on datasets like Transactions and MIMIC-IV. Additionally, we provide the first open-source CUDA implementation of the batched linear assignment solver, significantly enhancing the applicability of the proposed T-mAP metric.
@@
A detailed description of the benchmark can be found in Appendix C, with further details on HoTPP’s performance optimizations outlined in Appendix C.5.
4.1. Datasets

### v4

@@
Abstract
Forecasting multiple future events within a given time horizon is essential for applications in finance, retail, social networks, and healthcare. Marked Temporal Point Processes (MTPP) provide a principled framework to model both the timing and labels of events. However, most existing research focuses on predicting only the next event, leaving long-horizon forecasting largely underexplored. To address this gap, we introduce HoTPP, the first benchmark specifically designed to rigorously evaluate long-horizon predictions. We identify shortcomings in widely used evaluation metrics, propose a theoretically grounded T-mAP metric, present strong statistical baselines, and offer efficient implementations of popular models. Our empirical results demonstrate that modern MTPP approaches often underperform simple statistical baselines. Furthermore, we analyze the diversity of predicted sequences and find that most methods exhibit mode collapse. Finally, we analyze the impact of autoregression and intensity-based losses on prediction quality, and outline promising directions for future research. The HoTPP source code, hyperparameters, and full evaluation results are available at https://github.com/ivan-chai/hotpp-benchmark.
1 Introduction
@@
Event sequences differ fundamentally from other data types. Unlike tabular data [29], ESs include timestamps and possess an inherent order. In contrast to time series data [16], ESs are characterized by irregular time intervals and additional data fields, such as event types. These differences necessitate the development of specialized models and evaluation practices.
@@
Irregularly spaced timestamps are commonly modeled via Temporal Point Processes (TPP) [24]. Assigning categorical labels to these events leads to the Marked Temporal Point Process (MTPP) formalism. Since most event-sequence data can be viewed as MTPPs augmented with additional fields [18], accurately modeling event times and labels remains a central challenge in the broader event-stream (ES) domain.
@@
In practice, a common question arises: what events will occur, and when, within a specific time horizon? Forecasting multiple future events presents unique challenges that differ from traditional next-event prediction tasks [34, 40]. For instance, autoregressive event sequence prediction involves applying the model to its own potentially erroneous predictions. However, the challenges of autoregressive prediction in the context of MTPP have not been thoroughly explored. Another difficulty lies in evaluation. Methods like Dynamic Time Warping (DTW) are typically unsuitable for event sequences due to strict ordering constraints [28]. Some works have applied Optimal Transport Distance (OTD), a variant of the Wasserstein distance, to compare sequences of predefined lengths [20, 34, 33]. However, the limitations of this metric have not been previously considered.
@@
Figure 1: T-mAP evaluation pipeline. Unlike previous approaches, T-mAP evaluates sequences of variable lengths within the prediction horizon. It further improves performance assessment by analyzing label distributions rather than relying on fixed predictions.
@@
In this work, we provide the first in-depth analysis of models and metrics for long-horizon event forecasting, establishing a rigorous evaluation framework and a baseline for MTPP studies. To this end, we present HoTPP, a novel benchmark illustrated in Figure 1. Our key contributions are summarized as follows:
  1. 1.
@@
    We introduce the first open-source benchmark specifically designed for long-horizon event forecasting. By considering datasets and methods spanning finance, social networks, healthcare, and recommender systems, HoTPP significantly broadens the diversity and scale of data compared to previous benchmarks [33, 18, 32]. Moreover, we provide optimized training and inference algorithms to support large-scale studies.
  2. 2.
@@
    We demonstrate that widely used evaluation metrics for MTPPs overlook critical aspects of model performance. In particular, they are sensitive to calibration and incorrectly count false positives and false negatives. To address these issues, we introduce Temporal mean Average Precision (T-mAP), a novel evaluation metric inspired by best practices in computer vision [17]. Furthermore, we address a theoretical gap in computer vision by proving the correctness of the T-mAP computation algorithm.
  3. 3.
@@ 2 Related Work
Marked Temporal Point Processes. MTPP is a stochastic process consisting of a sequence of pairs $(t_{1},l_{1}),(t_{2},l_{2}),\dots$, where $t_{1}<t_{2}<\dots$ represent event times and $l_{i}\in\{1,\dots,L\}$ denote event type labels [24]. Common approaches to MTPP modeling focus on predicting the next event. A basic solution involves independently predicting the event time and type. A more advanced approach splits the original sequence into subsequences, one for each event type, and independently models each subsequence’s timing. Depending on the form of time-step distribution, the process is called Poisson or Hawkes.
@@
Over the last decade, the focus has shifted toward increasing model flexibility by applying neural architectures. Several works have employed classical RNNs [7, 31, 21] and transformers [41, 37], while others have proposed architectures with continuous time [19, 25, 14]. In this work, we evaluate MTPP models with both discrete and continuous time architectures. Unlike previous benchmarks, we also assess simple rule-based baselines and popular methods from related fields, including GPT-like prediction models for event sequences [18, 22]. For more details on MTPP modeling, refer to A.
@@
MTPP Evaluation. Previous benchmarks have primarily focused on medical data or traditional MTPP datasets. EventStream-GPT [18] and TemporAI [26] consider only medical data and do not implement methods from the MTPP field, despite their applicability. Early MTPP benchmarks, such as Tick [2] and PyHawkes [^1] https://github.com/slinderman/pyhawkes, implement classical machine learning approaches but exclude modern neural networks. While PoPPy [32] and EasyTPP [33] include neural methods, they do not consider rule-based and horizon prediction approaches, such as HYPRO [34] and Diffusion [40]. Furthermore, PoPPy does not evaluate long-horizon predictions at all. EasyTPP addresses this limitation to some extent by reporting the OTD metric, though it does not provide the corresponding evaluation code.
@@
Previous work also highlighted the limitations of Dynamic Time Warping (DTW) for event sequence evaluation [28], showing that DTW’s strict ordering constraints are impractical and should be avoided. This work proposed an alternative, the Order-preserving Wasserstein Distance (OPW). In our study, we address the limitations of OTD, a variant of the Wasserstein Distance, and propose a new metric called T-mAP. Unlike OPW, T-mAP operates on timestamps rather than event indices and involves two hyperparameters, compared to three in OPW.
3 T-mAP Metric
@@
In this section, we show that widely used evaluation metrics for long-horizon prediction often miscount false positives (FP) and false negatives (FN). We then introduce Temporal mean Average Precision (T-mAP), a novel, theoretically grounded metric that addresses these issues. In addition, T-mAP is invariant to linear calibration of model outputs, simplifying and stabilizing comparisons across different methods. A general overview of the T-mAP evaluation pipeline is presented in Fig. 1.
3.1 Limitations of the Next-Event and OTD Metrics
@@
Figure 2: Comparison of the OTD and T-mAP metrics. In the example 1.a, OTD’s prefix-based evaluation of three events erroneously produces a false negative, whereas T-mAP in example 1.b compares all events within the specified time horizon for a more accurate assessment. Moreover, OTD evaluates only the label with the highest probability, treating cases 2.a and 2.b as equivalent. By contrast, T-mAP considers the entire distribution, yielding a low score (0.33) for 2.a and a perfect score (1) for 2.b.
Recent studies have advanced the measurement of horizon prediction quality by employing OTD [20]. This metric is analogous to the edit distance between the predicted event sequence and the ground truth. Suppose there is a predicted sequence $S_{p}=\{(t^{p}_{i},l^{p}_{i})\}_{i=1}^{n_{p}},0<t_{1}^{p}<t_{2}^{p}<\dots$, and a ground truth sequence $S_{gt}=\{(t^{gt}_{j},l^{gt}_{j})\}_{j=1}^{n_{gt}},0<t_{1}^{gt}<t_{2}^{gt}<\dots$. These two sequences form a bipartite graph $\mathcal{G}(S_{p},S_{gt})$. The prediction is connected to the ground truth event if their types are equal. Denote $M(S_{p},S_{gt})$ a set of all possible matches in the graph $\mathcal{G}(S_{p},S_{gt})$. OTD finds the minimum cost among all possible matchings:
$\mathrm{OTD}(S_{p},S_{gt})=\min\limits_{m\in M(S_{p},S_{gt})}\left[\sum\limits_{(i,j)\in m}|t^{p}_{i}-t^{gt}_{j}|+C_{del}U_{p}(m)+C_{ins}U_{gt}(m)\right],$ (1)
where $U_{p}(m)$ is the number of unmatched predictions in matching $m$, $U_{gt}(m)$ is the number of unmatched ground truth events, $C_{ins}$ is an insertion cost, and $C_{del}$ is a deletion cost. It is common to take $C_{ins}=C_{del}$ [20]. It has been proven that OTD is a valid metric, as it is symmetric, equals zero for identical sequences, and satisfies the triangle inequality.
@@
3.2 Temporal mAP: Definition and Computation
In this section, we introduce a novel metric, Temporal mAP (T-mAP), which analyses all errors within a predefined time horizon, explicitly controls ordering, and is invariant to linear calibration. T-mAP is inspired by object detection methods from computer vision [8]. Object detection aims to localize objects within an image and identify their types. In event sequences, we tackle a similar problem but consider the time dimension instead of horizontal and vertical axes. Unlike object detection, where objects have spatial size, each event in an MTTP is a point without a duration. Therefore, we replace the intersection-over-union (IoU) similarity between bounding boxes with the absolute time difference.
@@ 3.2.1 Definition
3.2.2 Computation
Finding the optimal matching independently for each threshold value $h$ is impractical; thus, we need a more efficient method to evaluate T-mAP. This section shows how to optimize T-mAP computation using an assignment problem solver, like the Jonker-Volgenant algorithm [12]. The resulting complexity of T-mAP computation is $\mathcal{O}(BN^{3})$, where $B$ is the number of evaluated sequences and $N=\max(n_{p},n_{gt})$ is the number of events within the horizon.
Figure 3: T-mAP computation pipeline.
@@ 3.2.2 Computation
$\forall h\forall m\in M(\mathcal{G})\exists m_{h}\in M(\mathcal{G}_{h}):m_{h}\subset m.$ (3)
According to this theorem, we can compute the matching for the prediction $S^{l}_{gt}$ and subsequently reuse it for all thresholds $h$ and subsequences $S^{l}_{>}(h)$ to construct a precision-recall curve for the entire dataset, as shown in Fig. 3.c. The proof of the theorem, the study of calibration dependency, and the complete algorithm for T-mAP evaluation are provided in B.
3.2.3 T-mAP Hyperparameters
@@
Figure 4: T-mAP dependency on the $\delta$ parameter. The dashed line indicates the selected value. The results for all 5 datasets are presented in C.
3.2.4 Additional T-mAP Studies
We analyze simple synthetic datasets to further investigate the properties of the proposed T-mAP metric. In D, we demonstrate that T-mAP more effectively distinguishes trivial baselines from statistical methods in scenarios characterized by highly irregular time steps. Additionally, the results in E show that, unlike OTD, T-mAP captures model performance on rare (long-tail) classes.
4 HoTPP Benchmark
@@
The HoTPP benchmark integrates data preprocessing, training, and evaluation in a single toolbox. Unlike previous benchmarks, HoTPP introduces the novel T-mAP metric for long-horizon prediction. HoTPP also differs from prior MTPP benchmarks by including simple rule-based baselines and long-horizon prediction methods that simultaneously predict multiple future events, such as HYPRO [34] and Diffusion [40]. The HoTPP benchmark is designed to focus on simplicity, extensibility, evaluation stability, reproducibility, and computational efficiency.
@@
Simplicity and Extensibility. The benchmark code is organized into a clear structure, separating the core library, dataset-specific scripts, and configuration files. New methods can be integrated at various levels, including the configuration file, model architecture, loss function, metric, and training module. Core components are reusable through the Hydra configuration library [35]. For example, HoTPP can apply the HYPRO rescoring method to backbone models such as IFTPP, RMTPP, NHP, or ODE. Furthermore, each method’s backbone architecture can be easily switched between RNNs and Transformers. Methods such as IFTPP, RMTPP, and Diffusion are compatible with most HuggingFace models [30].
@@
Evaluation Stability. In the MTPP domain, many datasets contain only a few thousand sequences for training and evaluation. For example, the StackOverflow test set includes 401 sequences, Retweet contains 1956, and Amazon has 1851 sequences for testing. This possesses a unique challenge in the long-horizon forecasting task. Previous long-horizon evaluation pipelines [34, 33] made predictions only at the end of each sequence, resulting in a limited number of measurements and a low evaluation stability. To address this limitation, we evaluate long-horizon predictions at multiple intermediate points, decreasing the variance of evaluation metrics.
@@
Reproducibility. The HoTPP benchmark ensures reproducibility at multiple levels. First, we use the PytorchLightning library [9] for training with a specified random seed and report multi-seed evaluation results. Second, data preprocessing is carefully designed to ensure datasets are constructed reproducibly. Finally, we specify the environment in a Dockerfile.
@@
Computational Efficiency. Some methods are particularly slow, especially during autoregressive inference on large datasets. Straightforward evaluation can take several hours on a single Nvidia V100 GPU. We optimized the training and inference pipelines in two ways to accelerate computation. First, we implemented an efficient RNN that reuses computations during parallel autoregressive inference from multiple starting positions. Second, we developed highly optimized versions of the thinning algorithm used for sampling in NHP and continuous-time neural architectures (NHP and ODE) [24], achieving up to a 4x performance improvement compared to the official implementations. These optimizations allowed us to conduct the first large-scale evaluations of algorithms such as NHP, AttNHP, ODE, and HYPRO on datasets like Transactions and MIMIC-IV. Additionally, we provide the first open-source CUDA implementation of the batched linear assignment solver, significantly enhancing the applicability of the proposed T-mAP metric.
@@
A detailed description of the benchmark can be found in F, with further details on HoTPP’s performance optimizations outlined in G.
Table 1: Datasets statistics
Dataset # Sequences # Events Mean Length Mean Horizon # Classes OTD T-mAP Length Steps Cost $\delta$ Horizon Transactions 50k 43.7M 875 9.0 203 5 1 2 7 MIMIC-IV 120k 2.4M 19.7 6.6 34 5 2 4 28 Retweet 23k 1.3M 56.4 14.7 3 10 15 30 180 Amazon 9k 403K 43.6 14.8 16 5 1 2 10 StackOverflow 2k 138K 64.2 12.0 22 10 1 2 10
4.1 Datasets